# Efficient dynamic-nerf based volumetric video coding with rate distortion optimization

### v1

@@ Abstract
Index Terms—  Volumetric Videos, Dynamic NeRF, Compression, End-to-end Optimization
1 Introduction
@@
Fig. 1: The rendering results of our method in comparison with ReRF.
@@
Although MLP is a compact and powerful 3D representation, its high computational complexity poses significant challenges for applications that require real-time processing, such as volumetric video compression. Consequently, some works [6, 7, 8, 9] have focused on the utilization of explicit features like 3D grids [6], hash tables [7] and tensors [8] to accelerate the rendering of NeRF. However, the utilization of explicit features also incurs additional storage consumption, which can be particularly catastrophic for dynamic 3D scenes. Many works [10, 11, 12, 13, 14] have successfully achieved explicit feature compression of NeRF without compromising rendering speed in static scenes. However, there have been fewer efforts in compressing explicit features for dynamic scenes, with ReRF [15] being a representative work in this area.
@@
2 Method
Fig. 2: Training pipeline for the proposed method. (a) Initially, we load the basis field at the previous time step from the decoded frame buffer. During training, we only update the coefficient field and the residual field. (b) During end-to-end optimization, we estimate the rate of the coefficient field and the residual field as loss, using simulated quantization during the forward pass.
2.1 Preliminaries
@@
2.2 Compact Dynamic NeRF Representation
@@
Table 1: The bitrate allocation of different components in DiF. The basis grids occupy the majority bitrate of DiF, causing the necessity of fine compression.
@@ 2.3 End-to-end Optimization
3 Experiments
3.1 Implementation Details
@@
Fig. 3: The performance comparison of ReRF and our method in both ReRF and Dna-rendering dataset. BD-rate: Dna-rendering train:-66.54%, Dna-rendering test:-28.71%, ReRF train:-45.50%, ReRF test:-33.24%.
Fig. 4: Comparing the multi-view rendering results of our approach and ReRF on the DNA-rendering sequence, under similar bitrate.
3.2 Experimental Results
@@
Table 2: Step-by-step Analysis. The 2nd row shows the results of adding dynamic modeling to the baseline, and the 3rd row shows the results of the full model.
@@ 3.3 Ablation Studies
Baseline. We consider DiFDiF [9] as the baseline for our method. However, while DiF is designed for static scenes, when extending it to dynamic scenes, we employ DiF to model each time step individually. Additionally, to ensure a fair comparison, we utilize the 7zip which is a lossless compression algorithm to compress the DiF models.
@@
The results of the ablation studies are listed in Table 2. Compared to the baseline, our dynamic modeling method achieves a compression ratio of approximately 15 times while only incurring a minimal loss in PSNR. Furthermore, introducing end-to-end optimization further compresses the representation of dynamic NeRF without significantly impacting PSNR. The results of the ablation studies demonstrate the effectiveness of our proposed methods.
4 Conclusion

### v2

@@ Abstract
Index Terms:
Volumetric Videos, Dynamic NeRF, Compression, End-to-end Optimization
I Introduction
@@
Figure 1: The rendering results of our method in comparison with ReRF.
@@
Although MLP is a compact and powerful 3D representation, its high computational complexity poses significant challenges for applications that require real-time processing, such as volumetric video compression. Consequently, some works [6, 7, 8, 9] have focused on the utilization of explicit features like 3D grids [6], hash tables [7] and tensors [8] to accelerate the rendering of NeRF. However, the utilization of explicit features also incurs additional storage consumption, which can be particularly catastrophic for dynamic 3D scenes. Many works [10, 11, 12, 13, 14, 15] have successfully achieved explicit feature compression of NeRF without compromising rendering speed in static scenes. Recently, many works [16, 17, 18, 19, 20, 21] have extended NeRF to dynamic scenarios. However, there have been fewer efforts in compressing explicit features for dynamic scenes, with ReRF [22] being a representative work in this area.
@@
II Method
Figure 2: Training pipeline for the proposed method. (a) Initially, we load the basis field at the previous time step from the decoded frame buffer. During training, we only update the coefficient field and the residual field. (b) During end-to-end optimization, we estimate the rate of the coefficient field and the residual field as loss, using simulated quantization during the forward pass.
II-A Preliminaries
@@
II-B Compact Dynamic NeRF Representation
@@
TABLE I: The bitrate allocation of different components in DiF. The basis grids occupy the majority bitrate of DiF, causing the necessity of fine compression.
@@ II-C End-to-end Optimization
III Experiments
III-A Implementation Details
@@
Figure 3: The performance comparison of ReRF and our method in both ReRF and Dna-rendering dataset. BD-rate: Dna-rendering train:-66.54%, Dna-rendering test:-28.71%, ReRF train:-45.50%, ReRF test:-33.24%.
Figure 4: Comparing the multi-view rendering results of our approach and ReRF on the DNA-rendering sequence, under similar bitrate.
III-B Experimental Results
@@
TABLE II: Step-by-step Analysis. The 2nd row shows the results of adding dynamic modeling to the baseline, and the 3rd row shows the results of the full model.
@@ III-C Ablation Studies
Baseline. We consider DiF [9] as the baseline for our method. However, while DiF is designed for static scenes, when extending it to dynamic scenes, we employ DiF to model each time step individually. Additionally, to ensure a fair comparison, we utilize the 7zip which is a lossless compression algorithm to compress the DiF models.
@@
The results of the ablation studies are listed in Table II. Compared to the baseline, our dynamic modeling method achieves a compression ratio of approximately 15 times while only incurring a minimal loss in PSNR. Furthermore, introducing end-to-end optimization further compresses the representation of dynamic NeRF without significantly impacting PSNR. The results of the ablation studies demonstrate the effectiveness of our proposed methods.
IV Conclusion
@@
Acknowledgment
This work was supported by National Key R&D Project of China(2019YFB1802701), National Natural Science Foundation of China(62102024), the Fundamental Research Funds for the Central Universities; in part by the 111 project under Grant B07022 and Sheitc No.150633; in part by the Shanghai Key Laboratory of Digital Media Processing and Transmissions.